elimination/ addition
a - 2b = 7
3a - 2b = 9
with steps too

Answer by solver15(27)

first step is multipy a - 2b= 7 by -3
-3[a - 2b] = 7
so your new equation will be
-3a + 6b = -21
then eliminte:-3a + 6b = -21 by the second equation
-3a + 6b = -21
3a - 2b = 9
so -3a and + 3a is now elimated
4b/4 = -12/4
b = -3 the after that plug in -3 to any equation
we can pug it in a -2(-3) = 7
a +6 = 7
the subtract 6 to both sides
a = 1
so your solution will be (1,-3)